OpenStudy (anonymous):

I have a question about maximum turns on a graph. In this problem: F(x)=2x^3-2x^2+7x-25 The maximum number of turns would be 2 right? Since the largest exponent is 3 and you subtract 1 to get the turns?

OpenStudy (anonymous):

You are correct in your logic and application. Remember though that this is the MAX and you may not actually get all of them.....

OpenStudy (anonymous):

@hullsnipe do you mean in general or in this problem?

OpenStudy (anonymous):

In general. It all depends on the multiplicity of the factors (i.e., if you have two of the same factor you will have less overall turning points).

OpenStudy (anonymous):

oh okay! Thanks for the explanation.

OpenStudy (anonymous):

@hullsnipe Also, would you be able to explain how to find focal width? One of the problems I'm supposed to find the focal width for is x^2=12y

OpenStudy (anonymous):

degree - 1

OpenStudy (anonymous):

so yes, max is 3-1 = 2

OpenStudy (anonymous):

Standard Equation x2 = 4py Opens up or down Focus (0,p) Directrix y = -p Axis y – axis Focal Length p Focal Width \[\left| 4p \right|\] So since you are in standard form the "4p" = 12 which is the Focal Width

OpenStudy (anonymous):

Oh. Wow. I thought it was way more complicated. It makes sense now. Also, thanks for putting the focal length, saved me from asking.